Calculate the period of simple pendulum in the following lengths.

a) 3.0 m

b) 6.0

c) 9.0

d) 12.0


1
Expert's answer
2021-01-27T20:32:42-0500

a)

"T=2\pi\sqrt{\dfrac{L}{g}}=2\pi\sqrt{\dfrac{3.0\ m}{9.8\ \dfrac{m}{s^2}}}=3.47\ s."


b)

"T=2\pi\sqrt{\dfrac{6.0\ m}{9.8\ \dfrac{m}{s^2}}}=4.92\ s."


c)

"T=2\pi\sqrt{\dfrac{9.0\ m}{9.8\ \dfrac{m}{s^2}}}=6.02\ s."


d)

"T=2\pi\sqrt{\dfrac{12.0\ m}{9.8\ \dfrac{m}{s^2}}}=6.95\ s."